Ali spot a large octopus while diving 40 meters deep in the pacific ocean sea. Given the density of the sea water and the atmosphere pressure is 1025 kg/m^3 and 101.325 kPa



a.) Calculate the pressure exerted on Ali's body.


b.) Calculate the absolute pressure exerted on Ali's body.

1
Expert's answer
2021-11-17T11:19:10-0500

a) The pressure is


"p=\rho gh=401800\text{ Pa}."

b) The pressure:


"p_a=p+p_0=401800+101325=503125\text{ Pa}."